Adriano bought 9 soft and hard sweets in total. Soft sweets cost R3 each and hard sweets cost R2 each.If Adriano spent a total of R22, then how many soft sweets did he buy? 


1
Expert's answer
2020-10-15T16:40:15-0400

Let the number of soft sweets be ...."s" and the hard sweets be..... "h"

Hence

"s+h =9"

"s=9-h" and "h=9-s"


And the cost of soft sweets will be (R3)"s" and for hard sweets be (R2)"h"


Hence

"R(3s + 2h) = R22"

"3s+2h=22"

Replace value of h to be (9-s)

"3s + 2(9-s)=22"

"3s+18-2s=22"

"3s-2s=22-18"

"s=4"

"h=9-4=5"

Hence the number of soft sweets is 4

and number of hard sweets is 5